5. South Park Realtys stock had a 10% gain last year. According to the Fed, inflation was 3% last year. Assuming you had invested in South Park Realtys stock, what was your real rate of return? 6. Rocking Amp Bull energy DNA corp just paid a dividend of $3.10 a share last year. They have been increasing the dividends at a constant rate of 2.5% every year and plan on continuing to do so. If the market expects the total return on the stock to be 8% this year, what is the dividend yield?
Answer #1
5) (1+ Real Rate) x (1 + Inflation Rate ) = (1+ Nominal rate)
(1+ Real Rate) x (1 + 3% ) = (1+ 10%)
1+Real Rate = 1.1/1.03
1+Real Rate = 1.067961
Real Rate = 1.067961 -1
Real Rate = 0.067961 Or
6.7961%
6) Dividend Yield = Dividend/ Price x 100
Dividend = 3.1
Price (P0) = Dividend 1/(Re-g)
= 3.1 x(1.025)/(0.08-0.025)
= 57.77273
Dividend Yield = 3.1/57.77 x 100
5.3659%